# Long-RVOS: A Comprehensive Benchmark for Long-term Referring Video Object Segmentation

### v1

@@ Abstract
Referring video object segmentation (RVOS) aims to identify, track and segment the objects in a video based on language descriptions, which has received great attention in recent years. However, existing datasets remain focus on short video clips within several seconds, with salient objects visible in most frames. To advance the task towards more practical scenarios, we introduce Long-RVOS, a large-scale benchmark for long-term referring video object segmentation. Long-RVOS contains 2,000+ videos of an average duration exceeding 60 seconds, covering a variety of objects that undergo occlusion, disappearance-reappearance and shot changing. The objects are manually annotated with three different types of descriptions to individually evaluate the understanding of static attributes, motion patterns and spatiotemporal relationships. Moreover, unlike previous benchmarks that rely solely on the per-frame spatial evaluation, we introduce two new metrics to assess the temporal and spatiotemporal consistency. We benchmark 6 state-of-the-art methods on Long-RVOS. The results show that current approaches struggle severely with the long-video challenges. To address this, we further propose ReferMo, a promising baseline method that integrates motion information to expand the temporal receptive field, and employs a local-to-global architecture to capture both short-term dynamics and long-term dependencies. Despite simplicity, ReferMo achieves significant improvements over current methods in long-term scenarios. We hope that Long-RVOS and our baseline can drive future RVOS research towards tackling more realistic and long-form videos.
@@ 1 Introduction
Another concern lies in the evaluation metrics. Existing RVOS benchmarks [36, 20, 11, 7] typically evaluate performance by simply averaging the frame-wise segmentation metrics (e.g., $\cal{J}\&\cal{F}$). However, in real-world videos, the target objects do not appear in every frame, due to occlusion and constrained camera views. Therefore, a robust RVOS model should exhibit a sound temporal consistency. This means it should not only accurately segment the target when it is present, but also be able to predict its absence by outputting an empty mask. However, this capability of temporal consistency can not be adequately reflected by current metrics.
To address these gaps, this work proposes Long-RVOS, a large-scale benchmark for long-term video object segmentation. Long-RVOS is the first minute-level dataset in RVOS field, designed to tackle various realistic long-video challenges such as frequent occlusion, disappearance-reappearance and shot changing, as shown in Figure˜1 and Figure˜2. Additionally, we introduce two new metrics for better evaluation of temporal consistency: $\mathrm{tIoU}$, which measures the temporal overlap between predicted and ground-truth mask sequences; and $\mathrm{vIoU}$, which further measures the spatiotemporal volume overlap between them. We benchmark 6 state-of-the-art (SOTA) methods on Long-RVOS. The results demonstrate that while notable progress has been achieved in existing short-term benchmarks, these SOTA models still significantly struggle in realistic long-term scenarios, in both frame-level segmentation and video-level temporal consistency.
@@
Our contributions are summarized as follows: (i) We build Long-RVOS, the first large-scale benchmark for long-term RVOS. In Long-RVOS, we provide explicit description types and introduce new metrics to enable more comprehensive evaluation. (ii) We benchmark 6 state-of-the-art RVOS approaches on Long-RVOS, and propose a promising baseline ReferMo to address the challenges in long-video scenarios. These contributions establish a foundation for developing more robust RVOS models to handle the realistic long-term videos.
@@
Dataset Year Videos Mean duration Total duration Mean frames Masks Objects Object classes Text Text type
@@ 2 Related Works
RVOS Benchmarks. Given an object description, RVOS aims to identify, tracking and segment the referring object throughout the video. This task was initially introduced by Gavrilyuk et al. [11] and Khoreva et al. [20] in 2018, and has gradually become a popular topic in vision-language understanding. Gavrilyuk et al. [11] built A2D-Sentences and JHMDB-Sentences datasets, which focus on distinguishing different actors in a video through the descriptions about appearance and actions. Khoreva et al. [20] built Ref-DAVIS17 [20], which covers more diverse object types. Later, Ref-Youtube-VOS [36] was developed to further expand the benchmark scale in this field. Recently, MeViS [7] was proposed to highlight the importance of motion understanding in RVOS task. Despite the efforts, these benchmarks remain limited to short video clips lasting only a few seconds, with target objects clearly visible in most frames. Besides, they also lack sufficient evaluation mechanisms to consider the models’ specific capabilities in various aspects.
RVOS Approaches. Recent RVOS approaches are mainly based on Transformer-based end-to-end architecture, represented by MTTR [2] and ReferFormer [44]. For an effective and consistent object identification across the frames, follow-up works [30, 39, 14, 15] focus on integrating more object-level temporal information. ReferDINO [25] further improves the object-level visual-language understanding by inheriting the object grounding capability of GroundingDINO [28]. Meanwhile, the recent emergence of segment anything models, i.e., SAM [21] and SAM2 [35], provides unique opportunity for downstream segmentation tasks. Some frontier studies [47, 26, 5, 50, 1] explore to incorporate SAM and SAM2 into RVOS approaches, achieving significant improvements on existing benchmarks. For example, VideoLISA [1] incorporates large language models with SAM for reasoning video segmentation. SAMWISE [5] integrate text prompts into SAM2 by inserting trainable adapters. While these models achieve great progress in current short-video benchmarks, their abilities and robustness in handling real-world long videos is still unclear.
@@
Figure 3: Overview of object categories and scenes in Long-RVOS.
3 Long-RVOS: A Comprehensive Benchmark for Long-term RVOS
@@ 3.1 Video Collection
Previous RVOS datasets [11, 20, 36, 7] were typically constructed by providing text annotations on their corresponding VOS datasets (e.g., DAVIS17 [34], YouTube-VOS-2019 [46] and MOSE [8]). However, the existing long-term VOS datasets like LVOS [16] are limited in scale (containing only 720 videos), and most videos feature only one object target. Therefore, in order to establish a large-scale and diverse RVOS benchmark, we bypass the existing VOS datasets and turn to integrate multi-source long video datasets. Specifically, we build Long-RVOS based on three long-video datasets: TAO [6], VidOR [37], and Ego-Exo4D [12]. Moreover, TAO is a federated dataset combining multiple sources like Charades [38], LaSOT [9], ArgoVerse [4], AVA [13], YFCC100M [41], BDD-100K [49], and HACS [53]. We select videos and objects based on the following criteria:
@@
(a) Frequency distribution of the Top-50 object categories.
(b) Distribution of video duration.
(c) Distribution of object duration.
(d) Number of objects per video.
(e) Wordcloud of descriptions.
(f) Distribution of description types.
(g) Description number per object.
@@
3.3 Dataset Statistics
A detailed comparison with five existing RVOS datasets is shown in Table˜1. Notably, Long-RVOS offers significantly longer video duration than existing datasets. In addition, it contains the largest number of object classes and mask annotations. The large scale of Long-RVOS supports comprehensive training and evaluation of RVOS models.
Diverse Objects and Scenes. Long-RVOS is constructed by integrating multiple sources of video datasets, achieving a wide variety of objects and scenes, as illustrated in Figure˜3. These sources include indoor videos from Charades [38], outdoor videos from LaSOT [9], movie scenes from AVA [13], egocentric videos from Ego-Exo4D [12], and more diverse videos from other datasets [37, 41, 53]. In total, Long-RVOS contains 163 object categories, significantly surpassing the existing RVOS datasets. As shown in Figure˜4 (a), while Long-RVOS primarily focuses on human instances (71.9%), it also covers a diverse range of animals (12.7%) and instruments (15.4%). In Figure˜4 (b)-(d), we present further statistics on the videos and objects in Long-RVOS. Notably, the object number of each video spans from 2 to 14, preventing over-reliance on the most salient object and highlighting text-guided segmentation. With such extensive diversity, Long-RVOS can serve a comprehensive benchmark for RVOS research, facilitating the development of more real-world applications.
Diverse Descriptions. In real-world applications, user queries are always unpredictable. They might refer to salient attributes or instantaneous actions. To enable more comprehensive evaluation of model capabilities, Long-RVOS introduces three distinct types of text descriptions — static, dynamic, and hybrid. By explicitly categorizing these types, Long-RVOS prevents evaluation bias toward specific attribute cues (e.g., color or position), ensuring a fair and robust assessment. We present the detailed statistics of text descriptions in Figure˜4 (c)-(g). Critically, Long-RVOS maintains a balanced distribution of text types, and the description number for each object can vary from 1 to 9. These properties encourage comprehensive learning of diverse object attributes. With its explicit type annotations and diverse object descriptions, Long-RVOS provides a comprehensive benchmark for training and evaluating RVOS models in more realistic scenarios.
@@ 3.4 Evaluation Metrics
Formally, let $\hat{M}_{t},M_{t}\in\{0,1\}^{H\times W}$ denote the predicted and ground-truth masks at $t$-th frame, respectively, where $t\in[1,T]$. The frame-index sets of non-empty masks are defined as $\hat{\mathcal{T}}=\{t\mid\|\hat{M}_{t}\|_{0}>0\}$ (for predictions) and $\mathcal{T}=\{t\mid\|M_{t}\|_{0}>0\}$ (for the ground-truth), where the $\ell_{0}$-norm $\|\cdot\|_{0}$ denotes the count of non-zero elements. Then, $\mathrm{tIoU}$ is obtained by computing their IoU:
@@ 4 ReferMo: A Baseline Approach
As illustrated in Figure˜5, ReferMo decomposes the video into a sequence of clips, each consisting of a high-resolution keyframe and subsequent low-resolution motion frames. Then, it perceives the static appearance and short-term motion within local video clip, and captures the object target in long-term context by integrating the cross-clip information. Critically, ReferMo only predicts target masks over the keyframes, and the masks on the remain frames are generated by a pretrained object tracker (e.g., SAM2 [35]). In this way, ReferMo achieves a trade-off between training costs and long-term understanding without processing a large number of high-resolution frames.
@@ 4.1 Video Decomposition
Typically, a long-term video is composed of multiple shots, and the video frames within each shot often show significant temporal redundancy. This redundancy can be efficiently described by motion information to reduce the frame-by-frame computations. Inspired by Video-LaVIT [18], we employ the MPEG-4 [23] compression technique to extract keyframe and motion information from the videos. More sophisticated (but expensive) keyframe selection strategies [45, 42] can also be explored, but they are not the primary focus of this work. In MPEG-4, a video is decomposed into multiple clips, where each clip consists of a keyframe $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$ and the motion vectors $\mathcal{M}\in\mathbb{R}^{T\times\frac{H}{16}\times\frac{W}{16}\times 2}$ of its subsequent $T$ frames. Unlike the dense optical flow, these motion vectors can be directly extracted during the compressed video decoding process, making them well-suited for processing large-scale, long-term videos. The details of motion extraction process are provided in the supplementary.
@@ 4.2 From Local Perception to Global Interaction
Figure 6: Overview of local perceiver.
Different from the previous RVOS methods [30, 48, 25] that perform vision-language fusion on each single frame, we introduce motion representations to enable clip-level vision-language fusion. For each video clip, as shown in Figure˜6, the local perceiver encodes the text, keyframe and motion information through three separate encoders, and then employs a multi-modal fuser to progressively aggregate these information for clip-level object extraction. By collecting the objects across different video clips, we perform global temporal interaction to enable consistent object prediction and long-term temporal understanding.
@@
where $W_{down}^{I},W_{down}^{M}\in\mathbb{R}^{d\times r}$ indicate the low-rank projectors that compress the features to a lower dimension $r$, and $W_{up}\in\mathbb{R}^{r\times d}$ is a projector to resort the dimension. $\sigma$ denotes Sigmoid function and $\odot$ denotes Hadamard product. $\gamma\in\mathbb{R}^{d}$ is a learnable vector to modulate the channel-wise weights.
Vision-Language Fusion. We use the dual cross-attention modules [24, 28] for deep vision-language fusion. Formally, given the clip-level vision features $F\in\mathbb{R}^{N\times d}$ and the language features $E\in\mathbb{R}^{L\times d}$, where $N$ and $L$ individually denote their token number, we derive the cross-modal enhanced vision features $\widetilde{F}$ and language features $\widetilde{E}$ as follows:
$\mathcal{A}=\frac{F\cdot E^{\top}}{\sqrt{d}},\qquad\widetilde{F}=\mathrm{Softmax}(\mathcal{A})\cdot E,\qquad\widetilde{E}=\mathrm{Softmax}(\mathcal{A}^{\top})\cdot F.$ (5)
@@
Global Interaction. To enable consistent object prediction and long-term temporal understanding, we collect the object features across video clips to perform global temporal interactions. Following ReferDINO [25], we use the Hungarian algorithm [22] to align the objects clip-by-clip. Then, we perform temporal self-attention over the aligned object features to achieve global modeling. For better modality alignment, we also infuse the language information $\widetilde{E}$ into the object features through a cross-attention layer. Finally, the interacted object features are output to the segmentation head for generating instance masks. Note that these masks are only predicted for the key frame within each clip, serving as object anchors for SAM2’s mask propagation in subsequent frames.

### v2

@@ Abstract
Referring video object segmentation (RVOS) aims to identify, track and segment the objects in a video based on language descriptions, which has received great attention in recent years. However, existing datasets remain focus on short video clips within several seconds, with salient objects visible in most frames. To advance the task towards more practical scenarios, we introduce Long-RVOS, a large-scale benchmark for long-term referring video object segmentation. Long-RVOS contains 2,000+ videos of an average duration exceeding 60 seconds, covering a variety of objects that undergo occlusion, disappearance-reappearance and shot changing. The objects are manually annotated with three different types of descriptions to individually evaluate the understanding of static attributes, motion patterns and spatiotemporal relationships. Moreover, unlike previous benchmarks that rely solely on the per-frame spatial evaluation, we introduce two new metrics to assess the temporal and spatiotemporal consistency. We benchmark 7 state-of-the-art methods on Long-RVOS. The results show that current approaches struggle severely with the long-video challenges. To address this, we further propose ReferMo, a promising baseline method that integrates motion information to expand the temporal receptive field, and employs a local-to-global architecture to capture both short-term dynamics and long-term dependencies. Despite simplicity, ReferMo achieves significant improvements over current methods in long-term scenarios. We hope that Long-RVOS and our baseline can drive future RVOS research towards tackling more realistic and long videos.
1 Introduction
@@
Another concern lies in the evaluation metrics. Existing RVOS benchmarks [38, 22, 13, 9] typically evaluate performance by simply averaging the frame-wise segmentation metrics (e.g., $\cal{J}\&\cal{F}$). However, in real-world videos, the target objects do not appear in every frame, due to occlusion and constrained camera views. Therefore, a robust RVOS model should exhibit a sound temporal consistency. This means it should not only segment the target when it is present, but also be able to predict its absence by outputting an empty mask. However, this capability of temporal consistency can not be adequately reflected by existing metrics.
To address these gaps, this work proposes Long-RVOS, a large-scale benchmark for long-term video object segmentation. Long-RVOS is the first minute-level dataset in RVOS field, designed to tackle various realistic long-video challenges such as frequent occlusion, disappearance-reappearance and shot changing, as shown in Figure 1 and Figure 2. Additionally, we introduce two new metrics for better evaluation of temporal consistency: $\mathrm{tIoU}$, which measures the temporal overlap between predicted and ground-truth mask sequences; and $\mathrm{vIoU}$, which further measures the spatiotemporal volume overlap between them. We benchmark 7 SOTA methods on Long-RVOS. The results demonstrate that while notable progress has been achieved in existing short-term benchmarks, these SOTA models still significantly struggle in realistic long-term scenarios, in both frame-level segmentation and video-level temporal consistency.
@@
To summarize, our contributions are two folds. 1) We build Long-RVOS, the first large-scale long-term RVOS benchmark. In Long-RVOS, we provide explicit description types and introduce new metrics to enable more comprehensive evaluation. 2) We benchmark 7 SOTA approaches on Long-RVOS, and propose a promising baseline ReferMo to address the challenges in long-video scenarios. These contributions establish a foundation for developing more robust RVOS models to handle realistic long videos.
@@
Dataset Year Videos
Mean
duration
Total
duration
Mean
frames
Masks
Objects
Object
classes
Text
Text
type
@@ 2 Related Works
RVOS Benchmarks. Given an object description, RVOS aims to identify, tracking and segment the referring object throughout the video. This task was initially introduced by Gavrilyuk et al. [13] and Khoreva et al. [22] in 2018, and has gradually become a popular topic in vision-language understanding. Gavrilyuk et al. [13] built A2D-Sentences and JHMDB-Sentences, which focus on distinguishing different actors in a video through the descriptions about appearance and actions. Khoreva et al. [22] built Ref-DAVIS17 [22], which covers more diverse object types. Later, Ref-Youtube-VOS [38] was developed to further expand the benchmark scale in this field. Recently, MeViS [9] was proposed to highlight the importance of motion understanding in RVOS task. Despite the efforts, these benchmarks remain limited to short video clips lasting only a few seconds, with target objects clearly visible in most frames. Besides, they also lack sufficient evaluation mechanisms to consider the models’ specific capabilities in various aspects.
RVOS Approaches. Recent methods are primarily based on Transformer architecture, represented by MTTR [3] and ReferFormer [48]. For a consistent object identification across the frames, follow-up works [32, 41, 16, 17] focus on integrating more object-level temporal information. ReferDINO [27] further improves the visual-language understanding by inheriting the object grounding capability of GroundingDINO [30]. Meanwhile, the recent emergence of segment anything models, i.e., SAM [23] and SAM2 [37], provides unique opportunity for downstream segmentation tasks. Some frontier studies [51, 28, 7, 1] explore to incorporate SAM and SAM2 into RVOS approaches. For example, VideoLISA [1] incorporates large language models with SAM for reasoning video segmentation. SAMWISE [7] integrate text prompts into SAM2 through trainable adapters. While these models achieve great progress in current short-video benchmarks, their abilities and robustness in handling real-world long videos is still unclear.
@@
Figure 3: Frequency distribution of the Top-50 object categories.
3 Long-RVOS
@@ 3.1 Video Collection
Previous RVOS datasets [13, 22, 38, 9] were typically constructed by providing text annotations on their corresponding VOS datasets (e.g., DAVIS17 [36], YouTube-VOS-2019 [50] and MOSE [10]). However, the existing long-term VOS datasets like LVOS [18] are limited in scale (only 720 videos), and most videos feature only one object target. Therefore, in order to establish a large-scale and diverse RVOS benchmark, we bypass the existing VOS datasets and turn to multi-source long video datasets. Specifically, we integrate three long-video datasets: TAO [8], VidOR [39], and Ego-Exo4D [14]. Additionally, TAO is a federated dataset combining multiple sources like Charades [40], LaSOT [11], ArgoVerse [5], AVA [15], YFCC100M [43], BDD-100K [53], and HACS [57]. These datasets typically provide bounding box annotations on sparse frames. Then, we select videos and objects based on the following criteria:
@@
Figure 4: Overview of objects and scenes in Long-RVOS.
@@
3.3 Dataset Statistics
A detailed comparison with five existing RVOS datasets is shown in Table 1. Notably, Long-RVOS offers significantly longer video duration than existing datasets. In addition, it contains the largest number of object classes and mask annotations. The large scale of Long-RVOS supports comprehensive training and evaluation of RVOS models.
Diverse Objects and Scenes. Long-RVOS is constructed by integrating multiple sources of video datasets, achieving a wide variety of objects and scenes, as illustrated in Figures 3 and 4. These sources include indoor videos from Charades [40], outdoor videos from LaSOT [11], movie scenes from AVA [15], egocentric videos from Ego-Exo4D [14], and more diverse videos from other datasets [39, 43, 57]. In total, Long-RVOS contains 163 object categories, significantly surpassing the existing RVOS datasets. While Long-RVOS primarily focuses on human instances (71.9%), it also covers a diverse range of animals (12.7%) and instruments (15.4%). In Figure 5, we present further statistics on the videos and objects in Long-RVOS. Notably, the object number of each video spans from 2 to 14, preventing over-reliance on the most salient object and highlighting text-guided segmentation. Such extensive visual diversity ensures that models are tested against a wide array of complex, real-world scenarios.
Table 2: Distribution of description types.
Type Static Dynamic Hybrid
Percentage 35.03% 32.45% 32.52%
(a) Distribution of video duration.
(b) Distribution of object duration.
(c) Number of objects per video.
(d) Description number per object.
@@ 3.3 Dataset Statistics
Diverse Descriptions. In real-world applications, user queries are always unpredictable. They might refer to salient attributes or instantaneous actions. To enable more comprehensive evaluation of model capabilities, Long-RVOS introduces three distinct types of text descriptions — Static, Dynamic, and Hybrid. By explicitly categorizing these types, Long-RVOS prevents evaluation bias toward specific attribute cues (e.g., colors or positions), ensuring a fair and robust assessment. As shown in Table 2, Long-RVOS maintains a balanced distribution among these three description types. In addition, Figure 5(d) illustrates that the description number for each object can vary from 1 to 9. These properties encourage comprehensive learning of diverse object attributes. We also present the wordcloud of Long-RVOS is in Figure 6. Together, the diversity in both visual content and textual descriptions establishes Long-RVOS as a truly comprehensive benchmark for the training and evaluation of long-form RVOS models.
@@ 3.4 Evaluation Metrics
Figure 6: Wordcloud of descriptions.
Formally, let $\hat{M}_{t},M_{t}\in\{0,1\}^{H\times W}$ denote the predicted and ground-truth masks at $t$-th frame, respectively, where $t\in[1,T]$. The frame-index sets of non-empty masks are defined as $\hat{\mathcal{T}}=\{t\mid\|\hat{M}_{t}\|_{0}>0\}$ (for predictions) and $\mathcal{T}=\{t\mid\|M_{t}\|_{0}>0\}$ (for the ground-truth), where the $\ell_{0}$-norm $\|\cdot\|_{0}$ denotes the count of non-zero elements. Then, $\mathrm{tIoU}$ is obtained by computing their IoU as follows:
@@ 4 ReferMo: A Baseline Approach
As illustrated in Figure 7, ReferMo decomposes the video into a sequence of clips, each consisting of a high-resolution keyframe and subsequent low-resolution motion frames. Then, it perceives the static appearance and short-term motion within local video clip, and captures the object target in long-term context by integrating the cross-clip information. Critically, ReferMo only predicts target masks over the keyframes, and the masks on the remain frames are generated by a pretrained mask tracker (e.g., SAM2 [37]). In this way, ReferMo achieves a trade-off between training costs and long-term understanding without processing a large number of high-resolution frames.
@@ 4.1 Video Decomposition
Typically, a long-term video is composed of multiple shots, and the video frames within each shot often show significant temporal redundancy. This redundancy can be efficiently described by motion information to reduce the frame-by-frame computations. Inspired by Video-LaVIT [20], we employ the MPEG-4 [25] compression technique to extract keyframe and motion information from the videos. More sophisticated (but expensive) keyframe selection strategies [49, 46] can also be explored, but they are not the primary focus of this work. In MPEG-4, a video is decomposed into multiple clips, where each clip consists of a keyframe $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$ and the motion vectors $\mathcal{M}\in\mathbb{R}^{T\times\frac{H}{16}\times\frac{W}{16}\times 2}$ of its subsequent $T$ frames. Unlike the dense optical flow, these motion vectors can be directly extracted during the compressed video decoding process, making them well-suited for processing large-scale, long-term videos. More details of motion extraction process are provided in the supplementary.
@@ 4.2 From Local Perception to Global Interaction
Different from the previous RVOS methods [32, 52, 27] that perform vision-language fusion on each single frame, we introduce motion representations to enable clip-level vision-language fusion. For each video clip, as shown in the right part of Figure 7, the local perceiver encodes the text, keyframe and motion information through three separate encoders, and then employs a multi-modal fuser to progressively aggregate these information for clip-level object extraction. By collecting the objects across video clips, we perform global temporal interaction to enable consistent object prediction and long-term temporal understanding.
@@
where $W_{down}^{I},W_{down}^{M}\in\mathbb{R}^{d\times r}$ indicate the low-rank projectors that compress the features to a lower dimension $r$, and $W_{up}\in\mathbb{R}^{r\times d}$ resorts the dimension. $\sigma$ denotes Sigmoid function and $\odot$ denotes Hadamard product. $\gamma\in\mathbb{R}^{d}$ is a learnable vector to modulate the channel-wise weights.
Vision-Language Fusion. We use dual cross-attention [26, 30] for deep vision-language fusion. Formally, given the clip-level vision features $F\in\mathbb{R}^{N\times d}$ and the language features $E\in\mathbb{R}^{L\times d}$, where $N$ and $L$ individually denote their token number, we derive the cross-modal enhanced vision features $\widetilde{F}$ and language features $\widetilde{E}$ as follows:
$\left\{\begin{aligned} &\widetilde{F}=\mathrm{Softmax}(\mathcal{A})\cdot E,\\ &\widetilde{E}=\mathrm{Softmax}(\mathcal{A}^{\top})\cdot F,\end{aligned}\right.\qquad\text{where }\mathcal{A}=\frac{FE^{\top}}{\sqrt{d}}.$ (5)
@@
Global Interaction. To enable consistent object prediction and long-term temporal understanding, we collect the object features across video clips to perform global temporal interactions. Following ReferDINO [27], we use the Hungarian algorithm [24] to align the objects clip-by-clip. Then, we perform temporal self-attention over the aligned object features to achieve global modeling. For better modality alignment, we also infuse the language information $\widetilde{E}$ into the object features through a cross-attention layer. Finally, the interacted object features are sent to the segmentation head for generating instance masks. Note that these masks are only predicted for the key frame within each clip, serving as object anchors for SAM2’s mask propagation in subsequent frames. More details are present in the supplementary.
Table 3: Comparison on Long-RVOS test set. FPS is estimated at 360P on Nvidia A6000 GPUs, excluding the video loading time.
Method Static Dynamic Hybrid Overall FPS
Without SAM / SAM2
SOC [32] NeurIPS’23 39.3 71.8 33.9 38.8 73.2 34.2 37.7 71.9 32.5 38.6 72.3 33.5 53.8
MUTR [52] AAAI’24 42.8 72.6 38.7 41.2 73.5 37.7 42.4 72.3 38.1 42.2 72.8 38.2 20.4
ReferDINO [27] ICCV’25 50.9 73.6 46.0 45.4 73.8 41.5 48.7 73.1 44.0 48.4 73.5 43.9 46.4
With SAM / SAM2
VideoLISA [1] NeurIPS’24 17.7 65.8 14.0 12.7 71.4 5.3 11.5 70.0 4.6 14.0 69.0 8.1 6.6
GLUS [28] CVPR’25 25.2 61.8 21.6 27.2 62.7 23.9 24.8 60.3 20.6 25.7 61.6 22.0 3.6
SAMWISE [7] CVPR’25 41.3 65.5 31.3 40.4 67.6 31.3 41.0 66.9 30.8 40.9 66.6 31.1 7.0
RGA3 [45] ICCV’25 22.1 59.8 16.9 23.4 61.0 19.0 22.2 59.2 16.7 22.5 60.0 17.5 8.7
ReferMo (Ours) 55.8 73.6 47.5 49.3 74.2 42.4 53.3 72.9 45.4 52.9 73.6 45.2 52.5